Please help me solve this.Calculate the surface area of a cylinder if its height is 60 cm and the circumference of its base is 25,12.
Answer by Cromlix(4381)

Hi there,
Circumference = pi * diameter
25.12 = pi * diameter
Diameter = 25.12/pi
Diameter = 8 cm
Radius = 4 cm.
Surface Area of a cylinder
= 2*pi*r*h + 2*pi*r^2 (* times)
= 2*pi*4*60 + 2*pi*4^2
= 1507.96 + 100.53
= 1608.49 cm^2
Hope this helps :-)